Suppose that each day the price of a stock moves up 1/8 th of a point with probability

1/3 and down 1/8 th of a point with probability 2/3. The price fluctuations are

independent each day.

(a) What is the probability that the stock has a price gain in 6 days?
Answer #1

probability that the stock has a price gain in 6 days =P(at most 2 down in 6 days)=P(X<=2)=P(X=0)+P(X=1)+P(X=2)

=6C0(2/3)0(1/3)6+6C1(2/3)1(1/3)5+6C2(2/3)2(1/3)4=0.100137